# The Impact of Prompt Programming on Function-Level Code Generation

### v1

@@ VI-A Lessons learned
L1: The differences in the results of prompt techniques are not dramatic: We carefully designed a full factorial experiment to evaluate not only prompt techniques but also combinations of them in a prompt. Our analyses revealed that, while there was an impact of some prompt techniques on the generated functions, the results for most of the prompt techniques were not that different. For example, the difference in the Pass@1 rates for the prompts with the highest and lowest rates is only around 10-12 percentage points (see Figure 5), and the effect sizes of the complexities are mostly small or insignificant (see Table IV). These insights align with other studies that evaluate prompt techniques on code summarization [18] and generation [22], where the performance results of different prompt techniques such as CoT, few-shot, self-collaboration, among others, are also comparable. In contrast, we see clearer differences in the performance results of some prompt techniques when using benchmarks for math-related tasks or general question-answering [25, 44]. We conclude that a strong emphasis on prompt programming is not necessary in the context of function-level code generation using current-generation models.
@@
L3: There is a trade-off between correctness and maintainability when choosing prompt techniques: Our analysis revealed contrasting results: prompts with few-shot examples or function signatures improved correctness but increased complexity and number of code smells, while prompts that employed persona, CoT or package had lower passing rates but significantly enhanced code maintainability (see Tables IV,V for complexity and Figure 10 for code smells). While previous research suggests that the use of a persona in the prompt does not improve the outcome [46] but can improve the personalization and user experience [26], we believe that this only applies to simple personas such as “software developer”. However, our results indicate that personas can be more beneficial when used as a way to induce additional quality requirements e.g., “software developer who writes clean and simple code”. Recent work has also shown that personas can be beneficial for code generation when used in more complex approaches such as self-collaboration where multiple personas (e.g., requirement engineer, software tester, and a developer) are used together to iteratively construct the code in a systematic way [24].
@@ VI-B Implications
I3: Different LLMs have different sensitivity levels to the prompt techniques We argue that a single prompt technique does not have the same impact on the different aspects (correctness, similarity, or quality) of the generated code across all LLMs. We have seen that the three LLMs demonstrated different sensitivity levels to the prompt. For instance, we saw how the similarity scores of Llama3-generated functions were significantly impacted by CoT, while it showed no effect for GPT-4o and Mistral (see Figure 9). The error types for Mistral in Figure 16 did not seem to be strongly impacted by prompt techniques as they did for other LLMs. Note that these differences in the models do not necessarily come from the model size and the number of parameters it was trained on, but rather the underlying architecture it uses (which aligns with findings by Wang et al. [18]). This implies that when a software company integrates an LLM into its processes and provides employee training, it should develop specific guidelines tailored to the LLM, including recommendations for prompt techniques that align with the model’s characteristics e.g., if an LLM returns simple functions in general, so prompt techniques that impact complexity may not be needed.
I4: Determining the purpose of the code generation is essential for the use of prompt programming Depending on whether the intended use of the LLM is to support human developers or to completely automate code generation, prompt programming has different significance. While the use of some prompt techniques has significantly minimized the number of errors and code smells, we saw that many of these issues can be easily fixed by human developers, arguably requiring less time and effort than applying an additional prompt technique. For instance, the absence of a signature in a prompt causes TypeErrors when the LLM misjudges the number of arguments (see Figure 7). This bug may be more easily fixed by the human rather than re-prompting the LLM with the correct signature. On the other hand, prompt programming can be more valuable when the purpose is to automate code generation and return correct and maintainable code without the need for human intervention, especially to apply simple modifications or refactoring actions.
@@ VI-C Threats to validity
External validity. The main threats to external validity in this study are associated with the prompt techniques, the LLMs and the benchmark we utilized. There are many possible prompt techniques that can potentially impact code generation, such as self-collaboration [31], tree-of-thought [50], or providing the whole class as context. However, we decided to select common prompt techniques that can be practically applied by a typical software developer in most code generation tasks. Another important question is whether the use of more powerful LLMs can result in different findings and eliminate the need for prompt programming. We used three current-generation LLMs during the study, including GPT-4o (200B parameters) and Llama3 (70B parameters). Our replication package [35] also includes results for older LLMs (GPT-3.5, Llama2), showing that the prompt techniques affecting code generation in this study similarly impact older models.
Internal validity. Regarding internal validity, we acknowledge that LLMs can be sensitive to format or structure of the prompt [48], or even the order of the few-shot examples [21]. We used a fixed order of the prompt techniques that we believe represents a natural sentence flow, and we ensured to use it consistently across all prompts. In addition, the manual creation of few-shot examples for our dataset may have introduced a degree of subjectivity. We therefore involve the first three authors in the process, allowing them to discuss possible examples and select the two most representative ones.
@@ VII Conclusion
A possible future extension of our research is to evaluate to what extent our findings generalize to other code generation tasks (e.g., line completion, program repair, or the generation of full applications). It is plausible that some of the prompt techniques that did not show a meaningful positive impact on correctness in our experiments (e.g., chain-of-thought) turn out to be more relevant if the generation task is more complex. Additionally, there are other quality metrics, such as performance or energy efficiency, which should be studied in future work — particularly given that recent work indicates that AI-generated code frequently exhibits performance regressions [51].

### v2

@@ VI-A Lessons learned
L1: The differences in the results of prompt techniques are not dramatic: We carefully designed a full factorial experiment to evaluate not only prompt techniques but also combinations of them in a prompt. Our analyses revealed that, while there was an impact of some prompt techniques on the generated functions, the results for most of the prompt techniques were not that different. For example, the difference in the Pass@1 rates for the prompts with the highest and lowest rates is only around 10 percentage points (see Figure 5), and the effect sizes of the complexities are mostly small or insignificant (see Table V). These insights align with other studies that evaluate prompt techniques on code summarization [18] and generation [22], where the performance results of different prompt techniques such as CoT, few-shot, self-collaboration, among others, are also comparable. In contrast, we see clearer differences in the performance results of some prompt techniques when using benchmarks for math-related tasks or general question-answering [26, 48]. We conclude that a strong emphasis on prompt programming is not necessary in the context of function-level code generation using current-generation models.
@@
L3: There is a trade-off between correctness and maintainability when choosing prompt techniques: Our analysis revealed contrasting results: prompts with few-shot examples or function signatures improved correctness but increased complexity and number of code smells, while prompts that employed persona, CoT or package had lower passing rates but significantly enhanced code maintainability (see Tables V,VI for complexity and Figure 10 for code smells). While previous research suggests that the use of a persona in the prompt does not improve the outcome [50] but can improve the personalization and user experience [27], we believe that this only applies to simple personas such as “software developer”. However, our results indicate that personas can be more beneficial when used as a way to induce additional quality requirements e.g., “software developer who writes clean and simple code”. Recent work has also shown that personas can be beneficial for code generation when used in more complex approaches such as self-collaboration where multiple personas (e.g., requirement engineer, software tester, and a developer) are used together to iteratively construct the code in a systematic way [24].
@@ VI-B Implications
I3: Different LLMs have different sensitivity levels to the prompt techniques We argue that a single prompt technique does not have the same impact on the different aspects (correctness, similarity, or quality) of the generated code across all LLMs. We have seen that the three LLMs demonstrated different sensitivity levels to the prompt. For instance, we saw how the similarity scores of Llama3-generated functions were significantly impacted by CoT, while it showed no effect for GPT-4o and Mistral (see Figure 9). The error types for Mistral did not seem to be strongly impacted by prompt techniques as they did for other LLMs. Note that these differences in the models do not necessarily come from the model size and the number of parameters it was trained on, but rather the underlying architecture it uses (which aligns with findings by Wang et al. [18]). This implies that when a software company integrates an LLM into its processes and provides employee training, it should develop specific guidelines tailored to the LLM, including recommendations for prompt techniques that align with the model’s characteristics e.g., if an LLM returns simple functions in general, so prompt techniques that impact complexity may not be needed.
I4: Determining the purpose of the code generation is essential for the use of prompt programming Depending on whether the intended use of the LLM is to support human developers or to completely automate code generation, prompt programming has different significance. We manually inspected 40 randomly selected failing functions from different code levels, each of which had passed with at least one other prompt, to understand what caused the failures. We saw that while the use of some prompt techniques has significantly minimized the number of errors and code smells, many of these issues can be easily fixed by human developers, arguably requiring less time and effort than re-prompting the LLM and applying an additional prompt technique. For instance, the absence of a signature in a prompt causes TypeErrors when the LLM misjudges the number of arguments, or misses that the function is a part of a class until the signature with a self parameter is provided. Moreover, prompting the LLM with few-shot examples reduced AssertionErrors mostly because the original prompt lacked clear specifications for edge cases and input/output formats, which could be picked up from the examples and result in passed tests (see Figure 7). On the other hand, prompt programming can be more valuable when the purpose is to automate code generation and return correct and maintainable code without the need for human intervention, especially to apply simple modifications or refactoring actions.
@@ VI-C Threats to validity
External validity. The main threats to external validity in this study are associated with the prompt techniques, the LLMs and the benchmark we utilized. There are many possible prompt techniques that can potentially impact code generation, such as self-collaboration [32], AceCoder [33], or providing the whole class as context. However, we decided to select common prompt techniques that can be practically applied by a typical software developer in most code generation tasks. Another important question is whether the use of more powerful LLMs can result in different findings and eliminate the need for prompt programming. We used three current-generation LLMs during the study, including GPT-4o (200B parameters) and Llama3 (70B parameters). Our replication package [39] also includes results for older LLMs (GPT-3.5, Llama2), showing that the prompt techniques affecting code generation in this study similarly impact older models.
Internal validity. Regarding internal validity, we acknowledge that LLMs can be sensitive to format or structure of the prompt [51], or even the order of the few-shot examples [21]. To address this, we manually refine the prompts and ensure that they have the same level of detail, for example, by removing examples that may be described in the original prompt to not impact the few-shot analysis. We also used a fixed order of the prompt techniques that we believe represents a natural sentence flow, and we ensured to use it consistently across all prompts. In addition, the manual creation of few-shot examples for our dataset may have introduced a degree of subjectivity. We therefore involve the first three authors in the process, allowing them to discuss possible examples and select the two most representative ones. Furthermore, to avoid failures due to incorrect function names, we replace the generated signature with the correct one before testing, though failures due to incorrect parameter names may still occur. Finally, since we used recent LLMs, they may have been trained on the same open-source GitHub code we used. To reduce this risk, we avoided code-focused models like Codex and CodeLlama, which are known to be trained specifically on GitHub data. We also checked for memorization by following the method from Schäfer et al. [53] using the maximum similarity metric. Similarity was found to produce more meaningful results when identifying memorization compared to other techniques that focus solely on code structure [54]. However, it may still miss other forms of memorization, particularly those involving structural overlap. We found that for all models, 85% of the generated functions had a maximum similarity score below 0.4, and none were higher than 0.7. This suggests that the models produced solutions based on understanding the input, not memorizing training data.
@@ VII Conclusion
A possible future extension of our research is to evaluate to what extent our findings generalize to other code generation tasks (e.g., line completion, program repair, or the generation of full applications) and other prompt techniques. It is plausible that some of the prompt techniques that did not show a meaningful positive impact on correctness in our experiments (e.g., chain-of-thought) turn out to be more relevant if the generation task is more complex. Additionally, there are other quality metrics, such as performance or energy efficiency, which should be studied in future work — particularly given that recent work indicates that AI-generated code frequently exhibits performance regressions [55].